# The Impact of Auxiliary Patient Data on Automated Chest X-Ray Report Generation and How to Incorporate It

### v1

@@ Abstract
This study investigates the integration of diverse patient data sources into multimodal language models for automated chest X-ray (CXR) report generation. Traditionally, CXR report generation relies solely on CXR images and limited radiology data, overlooking valuable information from patient health records, particularly from emergency departments. Utilising the MIMIC-CXR and MIMIC-IV-ED datasets, we incorporate detailed patient information such as aperiodic vital signs, medications, and clinical history to enhance diagnostic accuracy. We introduce a novel approach to transform these heterogeneous data sources into embeddings that prompt a multimodal language model, significantly enhancing the diagnostic accuracy of generated radiology reports. Our comprehensive evaluation demonstrates the benefits of using a broader set of patient data, underscoring the potential for enhanced diagnostic capabilities and better patient outcomes through the integration of multimodal data in CXR report generation.
@@ 1 Introduction
Figure 1: The patient data from MIMIC-IV-ED associated with a CXR exam from MIMIC-CXR. This includes the exam’s images, the corresponding radiology report, and the associated image metadata. The findings and impression sections of the radiology report form the ground truth for CXR report generation. ED-specific data, such as medicine reconciliation and aperiodic vital signs, is also available for the patient.
@@
Chest X-ray (CXR) exams, which consist of multiple images captured during an imaging session, are essential for diagnosing and managing a wide range of conditions, playing a significant role in patient care. Radiologists interpret these exams and produce a written report with their findings. However, prompt reporting is hindered by a multitude of issues, including high patient volumes and limited availability of radiologists (Bailey et al., 2022).
Machine learning for automated CXR report generation is a promising solution that has garnered significant attention in the literature (Jones et al., 2021). By leveraging multimodal language models, exams can be rapidly interpreted and reported, potentially providing quick and reliable diagnostic insights crucial for decision-making, such as triaging patients. Models are often trained to generate the findings and impression sections of a radiology report (Figure 1), where the former details the interpretation of a patient’s exam and the latter summarises the most important findings. Potential benefits include enhanced radiologist effectiveness, a reduced workload, alleviation of the burden of report writing, and improved patient outcomes (Shen, 2021; Irmici et al., 2023).
@@
Early methods for CXR report generation produced a separate report for each image within an exam Wang et al. (2018). Later methods improved on this by considering all images of an exam to generate a single report Miura et al. (2021); Nicolson et al. (2024a), and incorporating prior exams for a patient Wu et al. (2022); Nicolson et al. (2024a). Additionally, including the reason for conducting the exam (the indication section in Figure 1) offered a further improvement Nguyen et al. (2023). This indicates that CXR report generation could benefit from the inclusion of a more comprehensive set of patient data.
@@
Modern patient record systems are another rich source of patient data, containing detailed information that may be valuable for CXR report generation. However, (1) the utility of this data has not been empirically investigated, and (2) it is unclear how to harmonise this heterogeneous data into a unified multimodal language model. This paper aims to address these two points. To achieve this, we combine CXR exams from MIMIC-CXR (Johnson et al., 2019) with emergency department (ED) patient records from MIMIC-IV-ED (Johnson et al., 2023). This means that for a single exam, a wide variety of multimodal data is available, as shown in Figure 1. From MIMIC-CXR, we utilise the images, their metadata, and several sections of the radiology report. Notably, incorporating the comparison or history section is a novel approach in the literature. From MIMIC-IV-ED, we investigate triage information, aperiodic vital signs, medications, and other data to provide a wider clinical context.
@@
We explore combining these sources of patient data as patient embeddings to prompt a multimodal language model. We demonstrate that complementary information from different data sources can improve the diagnostic accuracy of CXR report generation. To achieve this, we develop methods to transform tabular and aperiodic time series data into embeddings that can be used alongside token and image embeddings. We evaluate our model on MIMIC-CXR exams with accompanying patient data from MIMIC-IV-ED, using metrics shown to closely correlate with radiologists’ assessments of reporting (Yu et al., 2023). The main contributions of this work are:
@@
  • An investigation into how patient data impacts CXR report generation, focusing on the effects of specific data sources, such as medications and vital signs.
  • An empirical evaluation demonstrating that using multiple patient data sources — from a patient’s CXR exams and their ED record — significantly improves diagnostic accuracy.
@@
  • Introducing methods to convert multimodal patient data into embeddings for a language model, including numerical, categorical, free text, temporal, and image data.
@@
  • A release of dataset splits based on MIMIC-CXR and MIMIC-IV-ED, linking patient exams with their associated ED records (available as a Hugging Face dataset). This, along with our code repository and Hugging Face checkpoint can be found at: https://anonymous.4open.science/r/anon-D83E, enabling others to experiment with new methods for multimodal patient data.
@@ 2 Background and Related Work
There is evidence to suggest that incorporating more patient data improves diagnostic accuracy in radiology reporting. Initial improvements came from using multiple images per exam, like EMNLI, which often includes complementary frontal and lateral views of the patient Miura et al. (2021); Gaber et al. (2005). Methods such as CXRMate enhance diagnostic accuracy by incorporating a patient’s prior exams to identify changes over time Nicolson et al. (2024a); Wu et al. (2022); Kelly (2012). Including the indication section of the radiology report to provide clinical context also provides an improvement Nguyen et al. (2023). This trend indicates that providing more comprehensive patient data improves diagnostic accuracy, which we investigate in this work.
ED records contain a myriad of data, including vital signs such as respiratory rate, temperature, and blood pressure, which can aid in the identification of various diseases. A high respiratory rate and low blood oxygen saturation are indicative of conditions that compromise pulmonary function, such as pulmonary embolism. Similarly, an elevated body temperature is suggestive of an infectious process, such as pneumonia or tuberculosis. Incorporating such data into a CXR report generator could help corroborate subtle radiographic signs typical of these infections. Our findings demonstrate that patient data from the ED can indeed enhance CXR report generation.
@@
Recent advancements in integrating multimodal patient data have enhanced diagnostic and predictive healthcare capabilities. A study showed that a Transformer encoder combining imaging and non-imaging data outperformed single-modality models, diagnosing up to 25 conditions with higher AUC scores Khader et al. (2023b). Similarly, the MeTra architecture, which integrates CXRs and clinical parameters, demonstrated superior performance in predicting ICU patient survival compared to using either CXRs or clinical data alone Khader et al. (2023a). ETHOS, using a zero-shot learning approach, outperformed single-modality models in predicting inpatient mortality, ICU length of stay, and readmission rates Renc et al. (2024). These studies highlight the importance of multimodal data for improved healthcare analytics. Our work demonstrates that incorporating a comprehensive set of multimodal patient data enhances CXR report generation.
Recent advancements in multi-task learning have significantly improved biomedical models by leveraging shared knowledge. Med-PaLM M, a generalist biomedical model, excels in multiple tasks including classification, question answering, visual question answering (VQA), report summarisation, report generation, and genomic variant calling, using diverse input modalities like images, text, and genomics. It often outperforms specialised models, demonstrating superior performance and generalisation (Tu et al., 2024).
@@
Similarly, MIMIC-CXR has been leveraged for multi-task learning with models like MedXChat, which integrates instruction-tuning and Stable Diffusion to perform CXR report generation, VQA, and report-to-CXR generation, outperforming other LLM multi-task learners (Yang et al., 2023). RaDialog, another LLM-based method, combines visual features and pathology findings to generate accurate radiology reports and support interactive tasks, significantly improving clinical efficacy. CXR-LLaVA, a multimodal LLM integrating a vision transformer with a language model, outperformed models like GPT-4 Vision and Gemini Pro Vision in CXR report generation Lee et al. (2024).
@@
Determining the state-of-the-art CXR report generation model can be challenging due to the unavailability of some models and the lack of comparison to recent methods. The 2024 Shared Task on Large-Scale Radiology Report Generation (RRG24) aimed to address this by benchmarking models on a common leaderboard. The winning model, CXRMate-RRG24 Nicolson et al. (2024b), a derivative of CXRMate, emerged as a strong contender for state-of-the-art. In this work, we compare our model to established models (e.g., EMNLI) and recent benchmarks (e.g., CXRMate-RRG24, CXRMate, CXR-LLaVA, MedXChat, and RaDialog). We ensure a fair comparison by using available code or obtaining generated reports directly from the authors. Our findings indicate our model produces significantly better results than these models.
@@ 3 Dataset
We construct a dataset of 46 1064610646\,10646 106 patients by linking individual patient information from two separate sources: (1) CXR exams from MIMIC-CXR and (2) emergency records from MIMIC-IV-ED. Thus we consider MIMIC-CXR exams that occurred during an ED stay from MIMIC-IV-ED. Both datasets are publicly available and originate from the Beth Israel Deaconess Medical Center in Boston, MA.
MIMIC-CXR was formed by first extracting patient identifiers for exams performed in the ED between 2011–2016, and then extracting all exams for this set of patients from all departments between 2011–2016. Each exam includes a semi-structured free-text radiology report (Figure 1) that describes the radiological findings of the images, written by a practising radiologist contemporaneously during routine clinical care. All images and reports were de-identified to protect privacy. Sections from the radiologist reports were extracted using a modification [^1] https://anonymous.4open.science/r/anon-D83E of the official text extraction tool [^2] https://github.com/MIT-LCP/mimic-cxr/tree/master/txt in order to obtain the findings, impression, indication, history, and comparison sections.
MIMIC-IV-ED consists of de-identified data from ED stays between 2011–2019. The data was converted into a denormalised relational database with six primary tables: ED stays, diagnosis, medicine reconciliation, medicine administration, triage, and aperiodic vital signs. We do not consider the diagnosis table in this work, as it indicates the outcome of a patient’s ED stay. The patients of MIMIC-CXR can be linked to MIMIC-IV-ED via an identifier, allowing an ED specific dataset to be formed.
Example tables for a patient’s exam are shown in Figure 1. The dataset was formed by extracting a patient’s exams whose times (formed by the ‘StudyDate’ and ‘StudyTime’ columns of the metadata table) occurred within the ‘intime’ and ‘outtime’ of one of their ED stays. [^3] Exam 59128861 was removed as it overlapped with two separate ED stays for the patient. Exams with either a missing findings or impression section were not considered. Using the official splits of MIMIC-CXR, this gave a train/validation/test split of 45 5274552745\,52745 527/343/236 patients, 76 3987639876\,39876 398/556/958 exams, and 151 818151818151\,818151 818/1 13711371\,1371 137/1 81218121\,8121 812 CXRs. Each of these exams had one ED stay and triage row; 53% had at least one medicine reconciliation row with up to 106 rows; 62% had at least one vital signs row with up to 69 rows; and 37% had at least one medication administration row with up to 52 rows. Exams had an indication section 66% of the time with a maximum of 75 words, a history section 34% of the time with a maximum of 74 words, and a comparison section 97% of the time with a maximum of 129 words. Only one exam had both an indication and a history section.
@@ 4 Methods
The patient data from MIMIC-CXR and MIMIC-IV-ED for an exam are transformed into embeddings, which are used to prompt a multimodal language model to generate the findings and impression sections of the radiology report, as illustrated in Figure 2. Additionally, ‘Source’ embeddings differentiate the source of the data (e.g., the ‘chief complaint’ column from the triage table, the indication section, etc.), and time delta embeddings represent the time difference between an event and the exam. Standard embeddings, such as position and token embeddings, are also included. The patient data embeddings originate from three main groups: the tables of MIMIC-IV-ED; the report, images, and metadata of the current exam from MIMIC-CXR; and the patient’s prior exams (also originating from MIMIC-CXR). The prior exam and image embeddings are described in Section A and Subsection C.2, respectively.
@@ 4.1 Time, Position, & Source Embeddings
The ED information from MIMIC-IV-ED is typically recorded as discrete events, such as medications administered or vital signs measured, each with a specific timestamp. Events that occur closer to the time of the patient’s exam are generally more relevant for diagnostic purposes. To capture this, a time delta is calculated by subtracting the time of an event from the time of the exam. The exam time originates from MIMIC-CXR’s metadata table (Figure 3), whereas most of the MIMIC-IV-ED tables have event times for each row. As shown in Figure 3, the time delta is first converted to hours and then mapped using $1/\sqrt{\bm{\Delta}+1}$, assigning higher weights to events that occurred closer to the exam. The mapped time deltas are then passed through a feedforward neural network (FNN) defined as $f(\bm{\Delta}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$, where $\bm{W}_{\text{1}}\in\mathbb{R}^{1,2048}$, $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $f(\cdot)$ is the sigmoid linear unit (SiLU) activation function (Hendrycks and Gimpel, 2016), and $H$ is the hidden size of the decoder. This process generates the time delta embeddings, which are subsequently added to the embeddings of their respective sources. As shown in Figure 2, time delta embeddings are only applied to the prompt. Patient data from the current exam, such as the images, have a time delta of zero, while data from prior exams have a positive time delta.
@@
The position embeddings are ordered by the time delta (Figure 3). This is due to the rotary position embeddings of the decoder; tokens that are closer together are given more importance. Hence, the smaller the time delta, the closer the embedding’s position is to the report token embeddings. Following Nicolson et al. (2024a), each unique patient data source is given its own source embedding. This includes the images, each report section, each table’s text column and value-category columns (described in the next section), and prior images and report sections.
4.2 Tabular Data

### v2

@@ Abstract
This study investigates the integration of diverse patient data sources into multimodal language models for automated chest X-ray (CXR) report generation. Traditionally, CXR report generation relies solely on CXR images and limited radiology data, overlooking valuable information from patient health records, particularly from emergency departments. Utilising the MIMIC-CXR and MIMIC-IV-ED datasets, we incorporate detailed patient information such as vital signs, medicines, and clinical history to enhance diagnostic accuracy. We introduce a novel approach to transform these heterogeneous data sources into embeddings that prompt a multimodal language model; this significantly enhances the diagnostic accuracy of generated radiology reports. Our comprehensive evaluation demonstrates the benefits of using a broader set of patient data, underscoring the potential for enhanced diagnostic capabilities and better patient outcomes through the integration of multimodal data in CXR report generation.
@@ 1 Introduction
Figure 1: The patient data from MIMIC-IV-ED associated with a CXR exam from MIMIC-CXR. This includes the exam’s images, the corresponding radiology report, and the associated image metadata. The findings and impression sections of the radiology report form the ground truth for CXR report generation. Emergency-specific data, such as reconciled medicines and aperiodic vital signs, are also available for the patient.
@@
Chest X-ray (CXR) exams, which consist of multiple images captured during an imaging session, are essential for diagnosing and managing a wide range of conditions, playing a significant role in patient care. Radiologists interpret these exams and produce a written report with their findings. However, timely reporting is hindered by a multitude of issues, including high patient volumes and limited availability of radiologists (Bailey et al., 2022).
@@
Automated CXR report generation using multimodal language models is a promising solution (Jones et al., 2021). Potential benefits include enhanced radiologist effectiveness, streamlining report writing, and improved patient outcomes (Shen, 2021; Irmici et al., 2023). Early methods produced a separate report for each image within an exam Wang et al. (2018). Later methods improved on this by considering all images of an exam to generate a single report Miura et al. (2021); Nicolson et al. (2024a), and incorporating prior exams for a patient Wu et al. (2022); Nicolson et al. (2024a). Including the reason for the exam (the indication section in Figure 1) offered a further improvement Nguyen et al. (2023). This indicates that CXR report generation benefits from the inclusion of a more comprehensive set of patient data.
@@
Incorporating clinical information, including electronic health record (EHR) data, enhanced the interpretation accuracy, clinical relevance, and reporting confidence of radiologists’ findings Castillo et al. (2021). A growing push to integrate EHR systems into radiology workflows highlights the potential for CXR report generation models to leverage patient data directly Geeslin and Gaskin (2016). In this study, we aim to empirically investigate if such data can also improve CXR report generation. To facilitate this, we combine CXR exams from MIMIC-CXR (Johnson et al., 2019) with emergency department (ED) patient records from MIMIC-IV-ED (Johnson et al., 2023). This provides a wide variety of multimodal data per exam, as shown in Figure 1. From MIMIC-CXR, we utilise the images, their metadata, and several sections of the radiology report. Notably, we incorporate the history or comparison section of the report, which has not been investigated previously. From MIMIC-IV-ED, we incorporate triage data, aperiodic vital signs, medicines, and other data to provide a wider clinical context.
@@
We also investigate how to harmonise these heterogeneous data into patient data embeddings to prompt a multimodal language model. In doing so, we develop methods to transform tabular and aperiodic time series data into embeddings that can be used alongside token and image embeddings. We evaluate our model using metrics shown to closely correlate with radiologists’ assessments of reporting (Yu et al., 2023). Through our evaluation, we demonstrate that complementary information from different data sources can improve the diagnostic accuracy of CXR report generation. The main contributions of this work are:
@@
  • An investigation demonstrating how integrating specific patient data sources, such as medicines and vital signs, enhances CXR report generation and improves diagnostic accuracy.
@@
  • Introducing methods to convert numerical, categorical, text, temporal, and image data into embeddings for a multimodal language model.
@@
  • A dataset linking MIMIC-CXR exams with MIMIC-IV-ED records, along with the code and Hugging Face model (available at: https://anonymous.4open.science/r/anon-D83E).
@@ 2 Background and Related Work
Incorporating more patient data has improved diagnostic accuracy in radiology reporting. Initial improvements came from using multiple images per exam, like EMNLI; CXR exams often include complementary frontal and lateral views of the patient Miura et al. (2021); Gaber et al. (2005). Methods such as CXRMate enhance diagnostic accuracy by incorporating a patient’s prior exams to identify changes over time Nicolson et al. (2024a); Wu et al. (2022); Kelly (2012); Bannur et al. (2023); Hou et al. (2023). Including the indication section of the radiology report to provide clinical context also provides an improvement Nguyen et al. (2023). Our investigation focuses on leveraging a more comprehensive set of patient data to improve diagnostic accuracy.
ED records contain a wide range of data, as shown in Figure 1. The reconciled medicines may include furosemide, a diuretic commonly prescribed for managing fluid overload, often associated with conditions such as pulmonary edema or congestive heart failure. Elevated blood pressure and an increased heart rate in a patient’s vital signs may correlate with findings such as cardiomegaly or vascular changes. Vital signs such as high temperature, elevated respiratory rate, and low oxygen saturation, along with chief complaints of cough and shortness of breath, may suggest pneumonia. Incorporating such data could complement imaging evidence and provide additional context to support better predictions. Our findings demonstrate that ED patient data can indeed improve CXR report generation.
@@
Recent advancements in integrating multimodal patient data have improved diagnostic and predictive healthcare tasks. A Transformer encoder combining imaging and non-imaging data outperformed single-modality models in diagnosing multiple conditions Khader et al. (2023b). Similarly, the MeTra architecture, integrating CXRs and clinical parameters, excelled in predicting ICU survival Khader et al. (2023a), and ETHOS, with zero-shot learning, surpassed single-modality models in predicting mortality, ICU length of stay, and readmission rate Renc et al. (2024). These studies underscore the value of multimodal data, and our work demonstrates its benefits for CXR report generation.
@@
Multi-task learning has enhanced biomedical models by leveraging shared knowledge across tasks. Med-PaLM M, a generalist biomedical model, excels in classification, question answering, VQA, report summarisation, report generation, and genomic variant calling, using diverse modalities like images, text, and genomics, often outperforming specialised models (Tu et al., 2024). Similarly, MIMIC-CXR has been utilised in multi-task learning with models like MedXChat, which integrates instruction-tuning and Stable Diffusion for tasks like CXR report generation, VQA, and report-to-CXR generation, surpassing other LLM multi-task learners (Yang et al., 2023). RaDialog combines visual features and pathology findings to generate accurate radiology reports and enable interactive tasks, improving clinical efficacy. CXR-LLaVA, a multimodal LLM, outperformed models such as GPT-4 Vision and Gemini Pro Vision in CXR report generation Lee et al. (2024).
@@
Determining the state-of-the-art in CXR report generation is challenging due to model unavailability and limited comparisons with recent methods. The 2024 Shared Task on Large-Scale Radiology Report Generation (RRG24) aimed to address this by benchmarking models on a common leaderboard. The winning model, CXRMate-RRG24 Nicolson et al. (2024b), a derivative of CXRMate, emerged as a strong contender for state-of-the-art. In this work, we compare our model to established models (e.g., EMNLI) and recent benchmarks (e.g., CXRMate-RRG24, CXRMate, CXR-LLaVA, MedXChat, and RaDialog). We ensure a fair comparison by using available code or obtaining generated reports directly from the authors. Our evaluation indicates that our model represents a statistically significant improvement over these.
@@ 3 Dataset
We construct a dataset of 46 1064610646\,10646 106 patients by linking individual patient information from two separate sources: (1) CXR exams from MIMIC-CXR and (2) emergency records from MIMIC-IV-ED. Thus, we consider MIMIC-CXR exams that occurred during an ED stay from MIMIC-IV-ED. Both datasets are publicly available and originate from the Beth Israel Deaconess Medical Center in Boston, MA.
MIMIC-CXR was formed by first extracting patient identifiers for exams performed in the ED between 2011–2016, and then extracting all exams for this set of patients from all departments between 2011–2016. Each exam includes a semi-structured free-text radiology report (Figure 1) written by a practising radiologist contemporaneously during routine clinical care. Models are often trained to generate the findings and impression sections of a radiology report, where the former details the interpretation of a patient’s exam and the latter summarises the most important findings. All images and reports were de-identified to protect privacy. Sections from the radiologist reports were extracted using a modification of the official text extraction tool in order to obtain the findings, impression, indication, history, and comparison sections. [^1] https://anonymous.4open.science/r/anon-D83E
MIMIC-IV-ED consists of de-identified data from ED stays between 2011–2019. The data was converted into a denormalised relational database with six primary tables: ED stays, diagnosis, reconciled medicines, administered medicines, triage, and aperiodic vital signs. We do not consider the diagnosis table in this work, as it indicates the outcome of a patient’s ED stay. The patients of MIMIC-CXR can be linked to MIMIC-IV-ED via an identifier, allowing an ED-specific dataset to be formed.
Example tables for a patient’s exam are shown in Figure 1. The dataset was formed by extracting patient exams with times (formed by the ‘StudyDate’ and ‘StudyTime’ columns of the metadata table) that occurred within the ‘intime’ and ‘outtime’ of one of the patient’s ED stays. [^2] Exam 59128861 was removed as it overlapped with two separate ED stays of a patient. Events during an ED stay that occurred after the exam were removed. Exams with either a missing findings or impression section were not considered. Using the official splits of MIMIC-CXR, this gave a train/validation/test split of 45 5274552745\,52745 527/343/236 patients, 76 3987639876\,39876 398/556/958 exams, and 151 818151818151\,818151 818/1 13711371\,1371 137/1 81218121\,8121 812 CXRs. Further details are provided in Appendix A.
@@ 4 Methods
We develop a novel approach to transform different sources of patient data from MIMIC-CXR and MIMIC-IV-ED into embeddings; these are then used to prompt a multimodal language model to generate the findings and impression sections of the radiology report, as illustrated in Figure 2. Each embedding of the prompt is the summation of a patient data embedding, a source embedding, a position embedding, and a time delta embedding. Source embeddings differentiate the source of the datum, for example, the ‘chief complaint’ column of the triage table, the indication section, or an image. A time delta embedding represents the time difference between an event and the exam. The patient data embeddings originate from three main groups: the tables of MIMIC-IV-ED; the report, images, and metadata of the current exam from MIMIC-CXR; and the patient’s prior exams (also originating from MIMIC-CXR). The prior exam and image embeddings are described in Section B and Subsection D.2, respectively.
@@ 4.1 Time, Position, & Source Embeddings
Events from MIMIC-IV-ED, e.g., administered medicines, are timestamped and are more relevant as they occur closer to the exam time Ben Abacha et al. (2023). Hence, time delta embeddings are used to indicate this to the model. The time delta is the event time subtracted from the exam time, converted to hours and mapped using $\bm{D}=1/\sqrt{\bm{\Delta}+1}$, emphasising recent events. These mapped time deltas are processed via a feedforward neural network (FNN), $f(\bm{D}\bm{W}_{\text{1}})\bm{W}_{\text{2}}$, where $\bm{W}_{\text{1}}\in\mathbb{R}^{1,2048}$, $\bm{W}_{\text{2}}\in\mathbb{R}^{2048,H}$, $f(\cdot)$ is the SiLU activation (Hendrycks and Gimpel, 2016), and $H$ is the decoder’s hidden size. As shown in Figure 2, these embeddings are applied only to the prompt. The exam time for current and prior exam data was used as the event time for the time delta calculation.
@@
The position embeddings are ordered by the time delta (Figure 3). This is due to the rotary position embeddings of the decoder; tokens that are closer together are given more importance. Hence, the smaller the time delta, the closer the patient data embedding’s position is to the report token embeddings. Following Nicolson et al. (2024a), each unique patient data source is given its own source embedding. This includes the images, each report section, each table’s text column and value-category columns (described in the next section), prior images, and prior report sections.
4.2 Patient Data Embeddings: Tabular Data